OpenStudy (anonymous):

whats a term sequence? like for the question... Give the 5th and 10th term for the sequence formula below. tn = 5, 10 5/4, 10/4 4/5, 2/5 4/5, 10/4

OpenStudy (anonymous):

A term sequence is something of the nature: 1 2 3 4 5 etc Where in this case each following number is the last one plus one. So formally: \[n_i = n_{i-1} + 1\]

OpenStudy (anonymous):

I dont think I can see your sequence, so I cant help you with an anwer. Try to do it yourself, if you still cant copy the sequence here! :)

OpenStudy (anonymous):

im so confused. lol. Whats the sequence im suppose to be copying?

OpenStudy (anonymous):

You see in your question, you go "The formula below: Tn = " and then just a big blank haha. So I dont know what the formula is.

OpenStudy (anonymous):

oh sorry about that here is the rest of the rest of the formula: tn = 4/n

OpenStudy (anonymous):

well look at the formula. n stands for number of the term. so if you want term number 5 and 10, you replace n by that. You get 4/5 and 4/10 Or in other words, 4/5 and 2/5

OpenStudy (anonymous):

but what if the equation is like this tn = (n – 3)2. And it still wants us to use 5 and 10. So i replace N with 5 and i have T5=(N-3)2. and now?

OpenStudy (anonymous):

you forgot to replace the other n! youll get T5 = (5-3)2 = 2*2 = 4

OpenStudy (anonymous):

OH!

OpenStudy (anonymous):

hope that helps :)

OpenStudy (anonymous):

\[t_n = \frac{4}{n}\] To find 5th and the 10th term, In the above value of tn , put n= 5 once and n = 10. \[t_5 = \frac{4}{5}\] And now put n = 10 \[t_{10} = \frac{4}{10} = \frac{2}{5}\]

OpenStudy (anonymous):

well soati, i got 9.8.... And waterineyes, a fraction isnt an answer choice...:)

OpenStudy (anonymous):

What do you mean, you got 9.8?

OpenStudy (anonymous):

What do you want to tell @razzy ??

OpenStudy (anonymous):

ok so. T5=(10-3)squared . so 10-3=7. 7 times 7 is 49. that divded by 5 is 9.8

OpenStudy (anonymous):

also waterineyes just wrote the exact same thing I did for the original question lol

OpenStudy (anonymous):

waterineyes, i just wanted to say that from the answer choices, what you said isnt there. and thats not what i got either...

OpenStudy (anonymous):

C is the option can't you see??

OpenStudy (anonymous):

Wait I dont get it. If you want T5 you just replace N by 5 and get (5-3)^2 But if you want T10 replace N by 10 and get back (10-3)^2 = 49. Why did you divide by 5?

OpenStudy (anonymous):

OH. I was solving it like this >>> T5=(10-3)squared.....i replaced both Ns.

OpenStudy (anonymous):

@soati , I just gave the answer in more understandable form...

OpenStudy (anonymous):

lol please. anyway, do you get it now razzy?

OpenStudy (anonymous):

waterineyes, we already got that answer... ok lemme try it again.

OpenStudy (anonymous):

YES. I GET IT NOW. Thank you soati. You were loads of help!

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

ok wait. rofl im confused again. the first question you helped me with was the TN=4/N. And the answer was 4/5 and 2/5. but the question i asked you second, is tn = (n – 3)2... And when i do this, i get the same answer we got in the first problem! And its not an answer for the 2nd.

OpenStudy (anonymous):

How so? What answers did you get? We just did both of them, and if Im not mistaken we got 4 for the 5th term and 49 for the 10th?

OpenStudy (anonymous):

Oh. nvmmm. rofl. i was confused because no one really made it official what out answer was, and we were talking about question 1 when you were trying to teach me

OpenStudy (anonymous):

ahah sorry. the question is getting pretty long yes. So for problem one, 4/5 and 2/5. Problem two, 4 and 49!

OpenStudy (anonymous):

OK. So were on the same page now. xD

OpenStudy (anonymous):

Let us try one more example @razzy .. Let us try for, tn = 4n + 5 NOTE: In tn, n is only the suffix it is nothing to do with calculation.. Only n that matters is n attached with digit 4.. So, tn = 4n + 5 Find the 20th and 40th term of this Arithmetic Progression...

OpenStudy (anonymous):

k. im trying it right now.

OpenStudy (anonymous):

Okay, take your time...